OpenStudy (mathmath333):

Ten tickets are numbered 1,2,3...10.Six tickets are selected at random one at a time with replacement. The probability that the largest number appearing on the selected ticket is 7 is

OpenStudy (linn99123):

LOL your answer is C

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{Ten tickets are numbered}\ 1,2,3\cdots \ 10.\ \text{Six tickets are } \hspace{.33em}\\~\\ & \normalsize \text{selected at random one at a time with replacement. The }\hspace{.33em}\\~\\ & \normalsize \text{probability that the largest number appearing on the selected }\hspace{.33em}\\~\\ & \normalsize \text{ticket is 7 is :}\hspace{.33em}\\~\\ & a). \dfrac{7^{6}-1}{10^{6}} \hspace{.33em}\\~\\ & b). \dfrac{7^{6}-6^{6}}{ 10^{6}} \hspace{.33em}\\~\\ & c). \dfrac{6^{6}}{ 10^{6}} \hspace{.33em}\\~\\ & d). \dfrac{7^{6}+6^{6}}{ 10^{6}} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (linn99123):

;-; whale then

OpenStudy (linn99123):

gimmiw a minute

OpenStudy (anonymous):

@mathmath333 be careful! @AlexandervonHumboldt2 once made a header of his tutorial 2+2? and his post was deleted. your can be as well.

OpenStudy (mathmath333):

i have freedom of speech

OpenStudy (linn99123):

i am so sorry i am a moron tho

OpenStudy (linn99123):

B IT IS B

OpenStudy (mathmath333):

how

OpenStudy (linn99123):

*falls over dead* idek i made undeadknight do it

OpenStudy (linn99123):

*sobs*

OpenStudy (linn99123):

i tried but now i must hang my self

ganeshie8 (ganeshie8):

We have \(7^6\) ways to choose 6 tickets using the digits : {1, 2, 3, 4, 5, 6, 7} Out of them, there will be \(6^6\) ways to choose 6 tickes using just the digits : {1, 2, 3, 4, 5, 6} Therefore, there are \(7^6-6^6\) ways to choose 6 digits such that \(7\) is guaranteed to exist in the selection and being the greatest.